OpenStudy (anonymous):

PLEASE HELP ILL GIVE A MEDAL!!! Indicate which property is illustrated in Step 1. Step 1 7 + (4 + 10) + 0 = 7 + (10 + 4) + 0 Step 2 = (7 + 10) + (4 + 0) Step 3 = (7 + 10) + 4 Step 4 = 17 + 4 A. associative property of addition B. associative property of multiplication C. commutative property of addition D. identity property of addition

OpenStudy (anonymous):

@AnswerMyQuestions @iGreen @Conqueror

OpenStudy (anonymous):

plz help

OpenStudy (igreen):

You can take B out because there is no multiplication.

OpenStudy (anonymous):

ok.. yea i guessed that..

OpenStudy (igreen):

Associative Property of Addition: \(a + (b + c) = (a + b) + c\) Commutative Property of Addition: \(a + b = b + a\) Identity Property of Addition: \(a + 0 = a\) So which one do you think it is? @brainbuzzed5646

OpenStudy (anonymous):

Associative?

OpenStudy (anonymous):

i hope im right

OpenStudy (igreen):

No, I don't think so.

OpenStudy (anonymous):

it was communitive

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

Commutative just switches it around, that's what it shows in the parenthesis: \(7 + (\color{red}{4 + 10}) + 0 = 7 + (\color{red}{10 + 4}) + 0\)